OpenStudy (anonymous):

Table 1 shows the marks obtained by 40 candidates in a test. Given that the median mark is 35.5, (a) find the value of x and y (b) state the modal class (c) calculate the standard deviation of the marks

OpenStudy (anonymous):



OpenStudy (anonymous):

@zzr0ck3r can you help me with this ?

OpenStudy (zzr0ck3r):

I forgot stats, sorry. Its been to many years.

OpenStudy (anonymous):

oh gosh, really? hmm, it;s alright.. thank you anyway :))

ganeshie8 (ganeshie8):

Since there are a total of 40 candidates, median = 20th + 21st / 2

OpenStudy (anonymous):

median = 20th + 21st / 2 I dont understand this

ganeshie8 (ganeshie8):

scratch that

ganeshie8 (ganeshie8):

4 + x + y + 10 + 8 = 40 x + y = 18

OpenStudy (anonymous):

Ok, I get the same too. but I make it x= 18 - y so I can use it later for substitute. .

OpenStudy (anonymous):

or I shouldn't did that at the first place ?

ganeshie8 (ganeshie8):

not sure, sorry...

ganeshie8 (ganeshie8):

@eliassaab

OpenStudy (anonymous):

@richyw Can you help me?

OpenStudy (richyw):

I'm actually going to sleep now. hope you figure it out!

OpenStudy (anonymous):

@eliassaab i want your help for this question ...

OpenStudy (cwrw238):

i'm struggling with this . i agree with whats stated already. I'm not sure if there's enough information given to solve this problem.

OpenStudy (anonymous):

@cwrw238 I already gave enough information for this question. I wrote exactly what was written in my book.